Suppose 735 mg of quercetin in 100 mL water were extracted with an equal volume of isopropyl ether. The layers were separated and the ether was evaporated. The residue was found to contain 585 mg of quercetin. What is the distribution constant K?


1
Expert's answer
2021-03-29T06:01:16-0400

q = (735-585)/735

= 150/735 = 0.20408

q = Vwater/(Vwater +KDVether)

KD =((V water/q)-Vwater))/Vether

KD =(100/0.20408) -100ml)/1000 ml

= 3.9